Bryan’s father is now twice as old as Bryan is.  His father was four times as old as him 16 years ago.  How old is his father now?


1
Expert's answer
2022-05-23T16:38:51-0400

x - Bryans age, y - fathers age.

2x-y=0

y-16-= 4(x-16)


2x-y=0

y= 4x-48


2x-4x+48=0


-2x= -48

x= 24

y= 48

Answer:48